Project

General

Profile

02/25/2014 Minutes

Attendance

Evan, Dominick, Michael, Gavin

Detector Conditions

New APDs installed on dcm 7-4, APD checkout proceeding.

Several stop/start runs - RC controller died several times.
Partition 3 not running due to DAQ investigations. Essentially trying to understand the DCM dropouts.

4 hour spill interruption where we lost beam events. Something disrupted the spill server, unclear what. Evan will discuss with DAQ experts. Occurs roughly ever month or so. Process thinks it's affect. Evan gets emailed when this occurs - last hour of log empty. Happened overnight and Evan fixed it when he woke up and noticed. Networking folks should notice. Evan has tested email notification and will broaden to nova-daq-support. Suggestion to modify the trigger scalar to change colour if not updating, shifter should have noticed numbers not updating. Shifter could have noticed this in nearline...

HardwareWatchList will be showcased at Friday's Data Quality meeting and then at 1pm at this meeting we shall make a stab at the first list.

DDT deployed on FarDet but not spinning processes yet. Check if the filenames also have the "_t*" that Michael uses to find files for processing.

New nearline page almost ready to go live. Final implementations required for an OnMon, Shifter, Slice button toggles. Importantly the shifter toggle which will gather a handful of useful plots shifters really should be looking at!
Follow up with Nitin and Kuldeep on the status of this.

Michael: Started one machine processing in safe environment "unofficially" so Hayes can work on webpage implementations.
Speaking of "flashing" indicators, OnMon usually just shows FEB HitRate so Mark and Michael toying with the idea of a shuffle feature that flicks between the plots. Mark will work on the "OnMon Shuffle" though this is low priority. Discussed the usefulness of this and argued for an on/off toggle. Not clear it should be the default.